OpenStudy (anonymous):

ln(x-2)+ln(2x-3)=2lnx Solve for x

OpenStudy (anonymous):

I divide both sides with ln?

OpenStudy (kc_kennylau):

Some formulas for you: \[\ln a+\ln b = \ln ab\\ n\ln a = \ln a^n\]

OpenStudy (anonymous):

log is a function if you have \(f(x+2)-f(2x-3)=2f(x)\) you cannot divide by \(f\) to get \[x+2-(2x-3)=2x\]

OpenStudy (anonymous):

so I should divide like f(x+2)/(2x-3)?

OpenStudy (anonymous):

=2ln(x)?

OpenStudy (anonymous):

you can start with \[\ln(x-2)+\ln(2x-3)=2\ln(x) \] and then rewrite it as \[\ln((x-2)(2x-3))=\ln(x^2)\] using two properties of the logarithm

OpenStudy (anonymous):

one is as @kc_kennylau wrote \[\log(A)+\log(B)=\log(AB)\] that gives you the left hand side the other is that \[n\log(x)=\log(x^n)\] which gives you the right hand side

OpenStudy (anonymous):

I have ln(2x^2-7x+6)=lnx^2 but I dont know what to progress from there.

OpenStudy (kc_kennylau):

Remove the ln from both sides

OpenStudy (anonymous):

then, since the log is a one to one function, if \[\ln((x-2)(2x-3))=\ln(x^2)\] then \[(x-2)(2x-3)=x^2\] and it is algebra from here on in

OpenStudy (anonymous):

Oh, after I divide both sides by ln, I subtract the x^2 and factor right?

OpenStudy (anonymous):

oh no you are not dividing by \(\ln\)

OpenStudy (anonymous):

the log is a one to one function that means if \(\ln(A)=\ln(B)\) then you know \(A=B\) that is not division in any way

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

but yes, subtract \(x^2\) set it equal to zero and solve

OpenStudy (anonymous):

so the answer I got is x=6 and 1 , is that correct?

OpenStudy (anonymous):

you should get \[x^2-7 x+6 = 0\] or \[(x-6)(x-1)=0\] but now you have to be very careful

OpenStudy (anonymous):

the solution to the quadratic is \(x=6\) or \(x=1\) but that is not the solution to your original equation, the one with the log

OpenStudy (anonymous):

sometimes there is a negative solution that doesn't fit into the log, but in this case, we have two positive so both would work right?

OpenStudy (anonymous):

that is the right idea, but the wrong answer 1 is not a solution, even though it is positive that is because you have \(\ln(x-2)\) so even though 1 is positive, if you put \(x=1\) you get \(\ln(1-2)=\ln(-1)\) which is impossible

OpenStudy (anonymous):

it is not a matter of the number being positive or negative, it is a matter of the actual input being positive or negative hope this is clear

OpenStudy (anonymous):

Okay, I understand now, thank you for taking time to teach me.

OpenStudy (anonymous):

yw